# Trapped-ion Laser Cooling in Structured Light Fields

### v2

@@ II Carrier-free GS EIT Cooling
Figure 3: (a) Level structure and fields relevant to EIT cooling of ${}^{40}\text{Ca}^{+}$. $\Delta_{p},\Delta_{c}$ are detunings of pump and cooling laser fields coupling $|S_{1/2},m_{j}=-1/2\rangle\leftrightarrow|P_{1/2},m_{j}=1/2\rangle$ and $|S_{1/2},m_{j}=1/2\rangle\leftrightarrow|P_{1/2},m_{j}=1/2\rangle$ respectively. $\Delta_{r}$ is the repumper’s detuning from $\text{D}_{3/2}$ and $P_{1/2}$ at $B=0$. (b) Excitation spectrum as a function of the cooling beam’s detuning, for fixed $\Delta_{p}$. Insets show choice of $\Delta_{c}$ and AC stark shift $\delta$ relative to the motional frequency $\omega_{m}$ to be cooled, for both RW and SW implementations. (c) Comparison of cooling limit over a range of motional frequencies between simulation results (points) and analytical calculation (solid line) for low-intensity $\Omega_{c}$ beam. $\Omega_{p}$ and $\Delta_{c}$ are optimized for $\omega_{m0}=2\pi\times 3$ MHz and $\Delta_{p}=5\Gamma$. Triangular and circular dots correspond to RW and SW pump beams, respectively. (d) Experimental schematics of carrier-free EIT cooling using integrated photonics. The first builds on the design for Doppler cooling with a similar integrated SW cooling beam and free-space RW pump beam. The second uses integrated photonics for all beams. A vertically emitted pump beam may achieve high circular polarization purity Massai et al. (2022) relative to a vertically oriented $B$-field. The cooling beam emitted in TM10 has a SW-like beam profile and comes from a shallow angle to satisfy the polarization requirements discussed below. Motional modes with normal mode vector projection along the cooling axis labeled would be cooled in either case.
@@
EIT-based cooling generally requires a pump beam (with Rabi frequency $\Omega_{p}$) inducing atomic excitation undergoing quantum interference with a weaker cooling beam ($\Omega_{c}$) in an effective three-level system (Fig. 3a). This results in a Fano profile in the cooling beam’s excitation spectrum, with an absorption null at $\Delta_{c}=\Delta_{p}$ and a nearby bright peak separated by the AC stark shift $\delta\equiv\frac{1}{2}\sqrt{\Omega_{p}^{2}+\Delta_{p}^{2}}-\Delta_{p}$ induced by the pump beam Morigi et al. (2000b); Roos et al. (2000), as shown in Fig. 3b. In the usual configuration with RW optical fields, $\Delta_{c}$ is tuned to the EIT null to suppress carrier excitation, and $\delta$ is matched to the frequency $\omega_{m}$ of the motional mode to be cooled, such that its RSB is preferentially driven (Fig. 3b). The high contrast between the RSB and the BSB excitation compared to that achieved with typical dipole transitions employed in Doppler cooling gives rise to cooling with a low final state number Morigi et al. (2000a); Roos et al. (2000); Lechner et al. (2016). However the nonzero BSB excitation amplitude limits final phonon number achievable, and the Fano profile is usually derived for weak probe beam, and fails to describe actual dynamics for large $\Omega_{c}$.
@@
In a carrier-free configuration, the ion can be positioned at a zero-intensity nodal point of the $\Omega_{c}$ beam. The internal state dressing fundamental to EIT relies on intensity of the pump beam $\Omega_{p}$, which can be delivered either as a RW or as a SW with anti-node at the ion location, with comparable performance. With the carrier transition nulled by the spatial coherence of the SW $\Omega_{c}$ field, carrier-free EIT utilizes the pump beam with adjusted parameters to null the BSB excitation amplitude. As shown in Fig. 3b, the cooling beam’s detuning is chosen such that the BSB and RSB lie at the EIT null and peak, respectively. That is, $\Delta_{c}=\Delta_{p}+\omega_{m0}$ and $\delta=2\omega_{m0}$. As in resolved sideband cooling, carrier-free EIT in the LDR then in principle drives strictly the RSB only, thus allowing extremely low final $\bar{n}_{\text{ss}}$ Zhang et al. (2012). Additionally, due to suppression of carrier coupling by the cooling beam, higher $\Omega_{c}$ can be applied to realize faster cooling as compared to the standard RW EIT. Additionally, carrier-free EIT can be expected to offer a larger bandwidth of motional mode frequencies cooled, due to the larger range of detunings from $\Delta_{c}$ for which high contrast between the RSB and BSB excitation probabilities obtains (see Fig. 3b). Thus as compared to RW-based EIT cooling, carrier-free EIT cooling is expected to enable GS cooling of multiple motional modes, simultaneously to lower final phonon numbers. Fig. 3c shows calculated cooling limit in an approximate three-level system for the same pump beam detuning $\Delta_{p}=5\Gamma$ and center motional frequency $\omega_{m0}=2\pi\times 3$ MHz, as a function of motional frequency $\omega_{m}$. Under the low cooling beam intensity limit, $\bar{n}_{ss}$ is independent of $\Omega_{c}$ and whether the pump beam has RW or SW field profile. indicating both the agreement between analytic treatment (lines, and see Appendix) and master-equation simulation (points), along with the significantly lower $\bar{n}_{ss}$ achievable over broader bandwidths using the carrier-free configuration.
@@
The analytical treatment described in the Appendix accurately describes cooling rate and limit for a three-level model for weak $\Omega_{c}$, as validated against master equation simulation in Fig. 3c. To describe the cooling behavior with large $\Omega_{c}$ near the saturation limit where the analytical treatment breaks down, we focus on master equation simulations of the relevant eight-level system relevant to ${}^{40}\text{Ca}^{+}$ at $B=10$ G and examine the achievable cooling rate, phonon number limit, and bandwidth. Theree different couplings are present in the system: a $\hat{\sigma}_{+}$-polarized pump beam $\Omega_{p}$ between $|S_{1/2},m_{j}=-1/2\rangle$ and $|P_{1/2},m_{j}=+1/2\rangle$, a $\hat{\pi}$-polarized cooling beam $\Omega_{c}$ between $|S_{1/2},m_{j}=+1/2\rangle$ and $|P_{1/2},m_{j}=+1/2\rangle$, and a resonant $\hat{\epsilon}_{p}$-polarized repumper $\Omega_{r}$ between the $D_{3/2}$ and $P_{1/2}$ states. Similar to Doppler cooling inside the LDR, the carrier’s relative strength is stronger for smaller $\eta$, suggesting larger possible advantage for low $\eta$. However, carrier-free EIT’s EF in cooling rate is not strictly proportional to $\eta_{0}^{-2}$ because the carrier transition in the RW EIT has also been nulled to first order. Motivated by potential integrated geometries for beam delivery, we also quantify the dependence of these enhancements on imperfect polarization.
@@
Two practical candidate beam geometries for carrier-nulled EIT cooling are shown in Fig. 3d, based either on SW or HG delivery of the cooling beam. Motivated by the full level structure relevant to ${}^{40}\text{Ca}^{+}{}$ (Fig. 3a), we consider geometries to best satisfy the polarization requirement of $\Omega_{c}$ and $\Omega_{p}$. As detailed in the Appendix, in either case, the motional modes cooled are those with some projection along the cooling beam’s field gradient; in contrast to RW-based cooling, the gradient of the $\Omega_{p}$ beam does not enter to leading order, and hence both the SW or HG direction can be tuned with respect to the trap axis to cool a combination of axial and radial modes.
@@
Figure 4: 8L EIT cooling scan of RW (left) and SW (right) on $\Omega_{c}$ and $\omega_{m}$. The first and second rows showcase the cooling rate and final state phonon number respectively. $\Omega_{p}=2\pi\times 52.2(36.4)$ MHz and $\Delta_{c}=5.14\Gamma(5\Gamma)$ for SW(RW) EIT as optimized beam parameters for $\Delta_{p}=5\Gamma$ and center $\omega_{m0}=2\pi\times 3$ MHz. Crosses label points of maximal $W_{c}$. Compared to RW, SW EIT cooling achieves higher cooling rates and lower limits for a larger bandwidth of motional frequency.
@@
Fig. 4 shows the simulated $W_{c}$ and $\bar{n}_{ss}$ as a function of the cooling beam’s Rabi frequency $\Omega_{c}$ and different motional frequencies $\omega_{m}$. The calculations assume a RW pump detuning $\Delta_{p}=5\Gamma$ and repumper amplitude $\Omega_{r}=2\pi\times 10$ MHz in both cases, with $\Delta_{c}$ and $\Omega_{p}$ taken as the optimal values for both RW and SW configurations to target optimal cooling of a motional mode frequency $\omega_{m0}=2\pi\times 3$ MHz. We find an overall improvement using carrier-free EIT in both cooling rate and final phonon number over a significantly expanded parameter range, cooling $\omega_{m}$ between $52\pi\times 1.5-$ 5 MHz to $\bar{n}_{ss}<10^{-1}$.
@@
Figure 5: (a) Time evolution of SW and RW EIT starting at $\langle n\rangle=4$ for $\omega_{m0}=2\pi\times 3$ MHz. Starting at $\bar{n}=4$, SW EIT cooling reaches $\bar{n}_{ss}=0.003$ in $34\,\mu s$ using optimal Rabi frequency $\Omega_{sw}=2\pi\times 42$ MHz for fast cooling rate, while RW EIT takes $75\,\mu s$ to achieve $\bar{n}_{ss}=0.017$ with optimal $\Omega_{rw}=2\pi\times 14$. (b) and (c) Cooling rate and limit for parameters chosen to optimally cool $\omega_{m0}=2\pi\times 3$ MHz as a function of the the pump (b) and cooling (c) beam’s polarization impurity, assumed evenly distributed in the two undesired polarization components in each case, showing the relative insensitivity in cooling beam impurity as compared to the pump. (d) $W_{c}$ and $\bar{n}_{ss}$ for a configuration based on first-order HG delivery of the pump beam (Fig. 3d), as a function of the cooling beam’s TM10 mode power impurity (modeled as fraction of power in the fundamental Gaussian instead of the HG10 mode).
@@
In Fig. 5a, we simulate a GS cooling instance with the optimal Rabi frequencies that correspond to the highest cooling rate at $\omega_{m0}=2\pi\times 3$ MHz in Fig. 4. Starting near the Dopper cooling limit $\bar{n}=4$, carrier-free EIT cools $\bar{n}_{ss}$ to $0.003$, $7$ times lower than that of the RW, within half of the time. Furthermore, operation at lower $\Omega_{c}$ than maximizes $W_{c}$ would allow substantially lower final $\bar{n}_{ss}$ (Fig. 4a) still. These enhancements in the cooling performance, especially in $\bar{n}_{ss}$ over a broader motional mode bandwidth than obtained by RW EIT could considerably alleviate the cooling time overhead by reducing or eliminating the need for further SBC.
@@
For integrated implementations enabling practical carrier-free EIT cooling, delivery of the highly pure circular polarization required to realize the effective three-level system presents a challenge Massai et al. (2022). To guide design of experimental configurations realizing these functions, we quantify polarization impurities’ effects on the predicted performance. We note that implementations at larger magnetic field with larger frequency selectivity can significantly alleviate the required polarization purity Marinelli (2020); however, below we focus on the more challenging polarization purities required for low-field (few G) regimes often used, assuming $B=10$ G in the simulations.
@@
As detailed below, optical access from integrated photonics and $B$-field arrangements impose constraints on design that generally allow practical delivery of one beam with near-perfect polarization purity but compromises in the other’s. We consider the effects of both pump and cooling beams’ polarization impurities separately. Fig. 5b shows cooling rates and limits as a function of the pump beam’s power impurity $\epsilon$, evenly distributed between the other two polarization directions such that the polarization vector can be written as $\sqrt{1-\epsilon}\hat{\sigma}_{+}+\sqrt{\epsilon/2}(\hat{\sigma}_{-}+\hat{\pi})$. Fig. 5c shows the same for the cooling beam’s power impurity. Cooling is clearly strongly sensitive to the strong $\hat{\sigma}_{+}$ pump beam’s impurity: for impurities larger than approximately 1% in relative intensity, the advantage of SW EIT cooling is significantly compromised. Relatively more tolerance holds for the weaker $\hat{\pi}$ cooling beam’s impurity.
@@
These polarization and mode purity requirements inform requirements for potential experimental configurations for implementation of this cooling scheme. While linear polarization purity can straightforwardly be achieved well below 0.1% with integrated photonic addressing Mehta and Ram (2017), $<1\%$ purities are challenging to achieve for circularly polarized emission and present a challenge for design He and Li (2014); Spektor et al. (2023). Fig. 3d gives two possible experimental schemes, both utilizing integrated delivery of the cooling beam but based either on free-space or integrated delivery of the $\hat{\sigma}_{+}$ beam. In the latter case, a vertical B-field is chosen given the likely advantage of a surface-normal-propagating $\hat{\sigma}_{+}$ beam with high purity from passively robust structures Massai et al. (2022); Natarajan et al. (2024); $\pi$-polarized cooling light then requires vertical B-field, which is approximately achieved via grating emission of a TM-polarized waveguide mode at a shallow emission angle propagating close to parallel with the trap chip. The ${\sim}10\%$ tolerance on the $\hat{\pi}$-beam polarization purity allows for emission angles for quasi-TM-polarized cooling beams of approximately $72.5^{\circ}$ from normal.
@@
Naturally, the advantage of carrier-free EIT cooling relies on the stable driving of the ions at the nodal points of the cooling beam. For the three-level system, ref. Zhang et al. (2012) shows that $\bar{n}_{ss}$ degrades by 3×3$\times$3 × for a positional deviation from the cooling beam’s node of $0.017\lambda$ with $\lambda$ the cooling beam’s wavelength. This corresponds to approximately $1\%$ of the SW’s peak anti-node intensity. Ion addressing in a phase-stable SW with a similar level of position accuracy has been demonstrated with integrated photonicsVasquez et al. (2023), indicating that the required positioning precision is practically achievable in present traps with integrated photonics. The bottom design in Fig. 3d utilizes fully integrated beams Massai et al. (2022), with cooling field delivered through TM10 mode. Fig. 5d shows that simulated $\bar{n}_{ss}$ in the eight-level system degrades by an amount consistent with above for 1% power impurity in the fundamental mode, agreeing with Zhang et al. (2012). This sensitivity sets a design goal for experiments to suppress the fundamental mode’s relative intensity to approximately 1% level to avoid unwanted scattering.
@@
Our results indicate significant possible gains in GS cooling via carrier-free EIT cooling, simultaneously in motional mode frequency bandwidth, final phonon number, and cooling rate. By potentially significantly reducing requirements for subsequent SBC, these techniques may alleviate a key bottleneck in trapped-ion system runtime. Polarization purity in pump beam delivery is quantified as a key challenge for implementation at low magnetic fields, along with mode purity requirements when utilizing HG mode drives.

### v3

@@ I Carrier-free Doppler cooling
Figure 3: (a) Level structure and fields relevant to EIT cooling of ${}^{40}\text{Ca}^{+}$. $\Delta_{p}$ and $\Delta_{c}$ are detunings of the pump and cooling laser fields that couple $|S_{1/2},m_{j}=-1/2\rangle\leftrightarrow|P_{1/2},m_{j}=1/2\rangle$ and $|S_{1/2},m_{j}=1/2\rangle\leftrightarrow|P_{1/2},m_{j}=1/2\rangle$ respectively. $\Delta_{r}$ is the repumper’s detuning from $D_{3/2}$ and $P_{1/2}$ at $B=0$. (b) Excitation spectrum for a three-level system as a function of the cooling beam’s detuning, for fixed $\Delta_{p}$. Insets show choice of $\Delta_{c}$ and AC stark shift $\delta$ relative to the motional frequency $\omega_{m}$ to be cooled, for both RW and SW implementations. (c) Comparison of cooling limit over a range of motional frequencies between simulation results (points) and analytical calculation (solid line) for low-intensity $\Omega_{c}$ beam. $\Omega_{p}$ and $\Delta_{c}$ are optimized for $\omega_{m0}=2\pi\times 3$ MHz and $\Delta_{p}=5\Gamma$. Triangular and circular dots correspond to RW and SW pump beams, respectively. (d) Experimental schematics of carrier-free EIT cooling using integrated photonics. The first builds on the design for Doppler cooling with a similar integrated SW cooling beam and free-space RW pump beam. The second uses integrated photonics for all beams. A vertically emitted pump beam may achieve high circular polarization purity [41] relative to a vertically oriented $B$-field. The cooling beam emitted in TM10 has a SW-like beam profile and comes from a shallow angle to satisfy the polarization requirements discussed below. Motional modes with normal mode vector projection along the cooling axis labeled would be cooled in either case.
Experimental implementation of carrier-free Doppler cooling may employ the configurations shown in Fig. 1c. Though for practical beam waists, the HG TE10’s effective $\eta$ is lower than in a RW, optimal sideband coupling can be recovered by increasing $\Omega_{0}$ inside the LD regime to hold $\eta\Omega_{0}$ constant. This allows cooling radial modes of multi-ion strings without requiring precise positioning of multiple ions along a series of SW fringes. Outside the LD regime, however, cooling through higher-order sideband excitations (with coupling strengths proportional to $\eta^{k}$ for a $k$th order sideband) contribute significantly to cooling, which are attenuated for lower driving beam LD parameter $\eta$ and fixed first-order coupling $\eta\Omega$. In contrast the heating through spontaneous decay on a $k$th order sideband is proportional to $\eta_{0}^{k}$ which is independent of the driving beam profile. As a result, the EF in cooling rate achievable with carrier-free cooling decreases outside the LD regime when using lower $\eta$ values. For example, in our eight-level system simulations we observe that the EF for SW cooling with $\eta=\eta_{0}/4$ compared to the best case RW with $\eta=\eta_{0}$ drops to nearly unity for $\bar{n}=50$.
Our results quantify the performance of carrier-free Doppler cooling beyond the LD regime previously treated. Though using ${}^{40}\text{Ca}^{+}$ in the simulation as an example, similar advantage would hold for other ion species and are more favorable for lower intrinsic $\eta_{0}$ [42]. While the advantage offered is modest in comparison to well within the LD regime, our results indicate that such configurations can still be applied for highly excited ions. Therefore, early cooling stages can share the same cooling beam profile with carrier-free ground-state cooling schemes, for which more substantial enhancements are predicted.
II Carrier-free ground-state EIT Cooling
@@
EIT-based cooling generally requires a pump beam (with Rabi frequency $\Omega_{p}$) inducing atomic excitation undergoing quantum interference with a weaker cooling beam ($\Omega_{c}$) in an effective three-level system (Fig. 3a). This results in a Fano profile in the cooling beam’s excitation spectrum, with an absorption null at $\Delta_{c}=\Delta_{p}$ and a nearby bright peak separated in frequency by the AC Stark shift $\delta\equiv\frac{1}{2}\sqrt{\Omega_{p}^{2}+\Delta_{p}^{2}}-\Delta_{p}$ induced by the pump beam [43, 14], as shown in Fig. 3b. In the usual configuration with RW optical fields, $\Delta_{c}$ is tuned to the EIT null to suppress carrier excitation, and $\delta$ is matched to the frequency $\omega_{m}$ of the motional mode to be cooled, such that its red sideband is preferentially driven (Fig. 3b). The high contrast between the red and blue sideband excitation compared to that achieved with typical dipole transitions employed in Doppler cooling gives rise to cooling with a low final state number [13, 14, 44]. However, the nonzero blue sideband excitation amplitude limits final phonon number achievable, and furthermore, the Fano profile is usually derived for weak probe beam, failing to accurately describe dynamics for large $\Omega_{c}$.
@@
In a carrier-free configuration, the ion can be positioned at a zero-intensity nodal point of the $\Omega_{c}$ beam. The internal state dressing fundamental to EIT relies on intensity of the pump beam $\Omega_{p}$, which can be delivered either as a RW or as a SW with anti-node at the ion location, with comparable performance. With the carrier transition nulled by the spatial coherence of the SW $\Omega_{c}$ field, carrier-free EIT utilizes the pump beam with adjusted parameters to null the blue sideband excitation amplitude. As shown in Fig. 3b, the cooling beam’s detuning is chosen such that the blue and red sideband lie at the EIT null and peak, respectively. That is, $\Delta_{c}=\Delta_{p}+\omega_{m0}$ and $\delta=2\omega_{m0}$. As in resolved sideband cooling, carrier-free EIT in the LD regime then in principle drives strictly the red sideband only, thus allowing extremely low final $\bar{n}_{ss}$ [26]. Additionally, due to suppression of carrier coupling by the cooling beam, higher $\Omega_{c}$ can be applied to realize faster cooling as compared to the standard RW EIT. Lastly, carrier-free EIT can be expected to offer a larger bandwidth of motional mode frequencies cooled, due to the larger range of detunings from $\Delta_{c}$ for which high contrast between the red and blue sideband excitation probabilities obtains (see Fig. 3b). Thus as compared to RW-based EIT cooling, carrier-free EIT cooling is expected to enable ground-state cooling of multiple motional modes, simultaneously to lower final phonon numbers. Fig. 3c shows calculated cooling limit in an approximate three-level system for the same pump beam detuning $\Delta_{p}=5\Gamma$ and center motional frequency $\omega_{m0}=2\pi\times 3$ MHz, as a function of motional frequency $\omega_{m}$. Under the low cooling beam intensity limit, final state phonon number $\bar{n}_{ss}$ is independent of cooling beam Rabi frequency and whether the pump beam has RW or SW field profile. The agreement between analytic treatment (lines, and see Appendix) and master-equation simulation (points) points to significantly lower $\bar{n}_{ss}$ achievable over broader bandwidths using the carrier-free configuration.
@@
To describe the cooling behavior with large $\Omega_{c}$ near the saturation limit where the analytical treatment breaks down, we perform master equation simulations of the eight-level electronic system relevant to ${}^{40}\text{Ca}^{+}$ at $B=10$ G and examine the achievable cooling rate, phonon number limit, and bandwidth. Three different couplings are present in the system: a $\hat{\sigma}_{+}$-polarized pump beam $\Omega_{p}$ between $|S_{1/2},m_{j}=-1/2\rangle$ and $|P_{1/2},m_{j}=+1/2\rangle$ with the LD parameter $\eta_{p}$, a $\hat{\pi}$-polarized cooling beam $\Omega_{c}$ between $|S_{1/2},m_{j}=+1/2\rangle$ and $|P_{1/2},m_{j}=+1/2\rangle$ with the LD parameter $\eta_{c}$, and a $\hat{\epsilon}_{p}$-polarized resonant repumper $\Omega_{r}$ between the $D_{3/2}$ and $P_{1/2}$, with $\hat{\epsilon}_{p}$ defined as in the previous section. Similar to Doppler cooling inside the LD regime, the carrier’s relative strength is stronger for smaller $\eta$, suggesting larger possible advantage for low $\eta$. However, carrier-free EIT’s EF in cooling rate is not strictly proportional to $\eta_{0}^{-2}$ because the carrier transition in the RW EIT has also been nulled to first order. Motivated by potential integrated geometries for beam delivery, we also quantify the dependence of these enhancements on imperfect polarization and beam delivery.
@@
Two practical candidate beam geometries for carrier-free EIT cooling are shown in Fig. 3d, based either on SW or HG delivery of the cooling beam. Motivated by the full level structure relevant to ${}^{40}\text{Ca}^{+}{}$ (Fig. 3a), we consider geometries that best satisfy the polarization requirement of $\Omega_{c}$ and $\Omega_{p}$. As detailed in the Appendix, in either case, the motional modes cooled are those with some projection along the cooling beam’s field gradient; in contrast to RW-based cooling where cooling rate is proportional to $|\eta_{c}-\eta_{p}|^{2}$, the gradient of the $\Omega_{p}$ beam does not enter to leading order, and hence both the SW or HG direction can be tuned with respect to the trap axis to cool a combination of axial and radial modes. In the following discussion, we assume the carrier-free cooling beam has spatial gradient along the motional axis such that its LD parameter $\eta_{c}$ is equal to $\eta_{0}$. For the RW configuration simulations, to maximize $|\eta_{c}-\eta_{p}|$ while accommodating the polarization requirements, we consider RW cooling and pump beam perpendicular to each other and $45^{\circ}$ from the motional axis such that $\eta_{c}=-\eta_{p}=\eta_{0}/\sqrt{2}$, the choice that maximizes the achievable RW cooling rate.
@@
Figure 4: EIT cooling performance as a function $\Omega_{c}$ and $\omega_{m}$. The top and bottom rows show the cooling rate and final state phonon number respectively, for typical RW (left) and carrier-free (right) configurations. For all plots, $\Delta_{p}=5\Gamma$, and we choose $\Omega_{p}=2\pi\times 52.2(36.4)$ MHz and $\Delta_{c}=5.14\Gamma(5\Gamma)$ for SW(RW) EIT as optimal parameters to cool a mode with $\omega_{m0}=2\pi\times 3$ MHz. Crosses label points of maximal $W_{c}$. Compared to RW, SW EIT cooling achieves higher cooling rates and lower limits for a larger bandwidth of motional frequency.
@@
Fig. 4 shows the simulated $W_{c}$ and $\bar{n}_{ss}$ as a function of the cooling beam’s Rabi frequency $\Omega_{c}$ and different motional frequencies $\omega_{m}$. The calculations assume a RW pump detuning $\Delta_{p}=5\Gamma$ and repumper amplitude $\Omega_{r}=2\pi\times 10$ MHz in both cases, with $\Delta_{c}$ and $\Omega_{p}$ taken as the optimal values for both RW and SW configurations to target optimal cooling of a motional mode frequency $\omega_{m0}=2\pi\times 3$ MHz. We find an overall improvement using carrier-free EIT in both cooling rate and final phonon number over a significantly expanded parameter range, cooling $\omega_{m}/2\pi$ between $1.5$ and $5$ MHz to $\bar{n}_{ss}<10^{-1}$.
@@
Figure 5: (a) Cooling trajectory for SW and RW EIT cooling. For a motional mode with frequency $\omega_{m}=2\pi\times 3$ MHz and starting at $\bar{n}=4$, SW EIT cooling reaches $\bar{n}_{ss}=0.003$ in $34\,\mu s$, while RW EIT takes $75\,\mu s$ to achieve $\bar{n}_{ss}=0.017$. In both cases, $\Omega_{c}$ is chosen to maximize $W_{c}$, corresponding to $\Omega_{c}=2\pi\times 42(14)$ MHz for the SW(RW) case. (b) and (c) Cooling rate and limit for parameters chosen to optimally cool $\omega_{m}=2\pi\times 3$ MHz as a function of the the pump (b) and cooling (c) beam’s polarization impurity, assumed evenly distributed in the two undesired polarization components in each case. The results indicate both the sensitivity of the carrier-free cooling’s enhancement to the pump beam purity, and the relative insensitivity to the cooling beam impurity. (d) $W_{c}$ and $\bar{n}_{ss}$ for a configuration based on first-order HG delivery of the pump beam (Fig. 3d), as a function of the cooling beam’s HG TM10 mode power impurity (modeled as fraction of power in the fundamental Gaussian instead of the HG10 mode).
@@
In Fig. 5a, we simulate a ground-state cooling instance with the optimal Rabi frequencies that correspond to the highest cooling rate at $\omega_{m0}=2\pi\times 3$ MHz in Fig. 4. Starting near the Dopper cooling limit $\bar{n}=4$, carrier-free EIT cools $\bar{n}_{ss}$ to $0.003$, $7$ times lower than that of the RW, within half of the time. Furthermore, operation at lower $\Omega_{c}$ than maximizes $W_{c}$ would allow substantially lower final $\bar{n}_{ss}$ (Fig. 4a) still. These enhancements in the cooling performance, especially in $\bar{n}_{ss}$ over a broader motional mode bandwidth than obtained by RW EIT could considerably alleviate the cooling time overhead by reducing or eliminating the need for further sideband cooling.
@@
For integrated implementations enabling practical carrier-free EIT cooling, delivery of the highly pure circular polarization required to realize the effective three-level system presents a challenge [41]. To guide design of experimental configurations realizing these functions, we quantify polarization impurities’ effects on the predicted performance. We note that implementations at larger magnetic field with larger frequency selectivity can significantly alleviate the required polarization purity [45]; however, below we focus on the more challenging polarization purities required for low-field (few G) regimes often used, still assuming $B=10$ G as above.
@@
Optical access from integrated photonics and $B$-field arrangements impose constraints on design that generally allow practical delivery of one beam with near-perfect polarization purity but compromises in the other’s. We consider the effects of both pump and cooling beams’ polarization impurities separately. Fig. 5b shows cooling rates and limits as a function of the pump beam’s power impurity $\epsilon$, evenly distributed between the other two polarization directions such that the polarization vector can be written as $\sqrt{1-\epsilon}\hat{\sigma}_{+}+\sqrt{\epsilon/2}(\hat{\sigma}_{-}+\hat{\pi})$. Fig. 5c shows the same for the cooling beam’s power impurity. Cooling is clearly sensitive to the strong $\hat{\sigma}_{+}$ pump beam’s impurity: for impurities larger than approximately 1% in relative intensity, the advantage of carrier-free EIT cooling is significantly compromised. Cooling performance is significantly more tolerant to the impurity of the $\hat{\pi}$-polarized cooling beam.
@@
These polarization and mode purity requirements inform potential experimental configurations for implementation of this cooling scheme. While linear polarization purity can straightforwardly be achieved well below 0.1% with integrated photonic addressing [46], achieving $<1\%$ circularly polarization purity presents a challenge for design [47, 48]. Fig. 3d gives two possible experimental schemes, both utilizing integrated delivery of the cooling beam but based either on free-space or integrated delivery of the $\hat{\sigma}_{+}$ beam. In the latter case, a vertical B-field is chosen given the likely advantage of a surface-normal-propagating $\hat{\sigma}_{+}$ beam with high purity from passively robust structures [41, 49]; $\pi$-polarized cooling light then corresponds to a vertically polarized $E$-field, which is approximately achieved via grating emission of a TM-polarized waveguide mode at a shallow emission angle propagating close to parallel with the trap chip. The ${\sim}10\%$ tolerance on the $\hat{\pi}$-beam polarization purity allows for emission angles for quasi-TM-polarized cooling beams of approximately $72.5^{\circ}$ from normal.
@@
Naturally, the advantage of carrier-free EIT cooling relies on the stable driving of the ions at the nodal points of the cooling beam. For the three-level system, ref. [26] shows that $\bar{n}_{ss}$ degrades by 3×3$\times$3 × for a positional deviation from the cooling beam’s node of $0.017\lambda$ with $\lambda$ the cooling beam’s wavelength. Consistent with results from our eight-level simulations (see Appendix), this corresponds to a position deviation from the SW node of order 10 nm, to a location with approximately $1\%$ of the SW’s anti-node intensity. Ion addressing in a phase-stable SW with a similar level of position accuracy has been demonstrated with integrated photonics [28], indicating that the required positioning precision is practically achievable. The bottom design in Fig. 3d utilizes fully integrated beams [41], with cooling field delivered through TM10 mode. Fig. 5d shows that simulated $\bar{n}_{ss}$ in the eight-level system degrades by an amount consistent with above for 1% power impurity in the fundamental mode, agreeing with [26]. This sensitivity sets a design goal for experiments to suppress the fundamental mode’s relative intensity to approximately 1% level to avoid unwanted scattering.
@@
Our results indicate significant possible gains in ground-state cooling via carrier-free EIT cooling, simultaneously in motional mode frequency bandwidth, final phonon number, and cooling rate. By potentially significantly reducing requirements for subsequent sideband cooling, these techniques may alleviate a key bottleneck in trapped-ion system runtime. Polarization purity in pump beam delivery is quantified as a key challenge for implementation at low magnetic fields, along with mode purity requirements when utilizing HG mode drives.
Integration of photonic elements for the UV wavelengths required for ${}^{40}\text{Ca}^{+}$ laser cooling presents a challenge but may leverage alumina or hafnia/alumina waveguides recently demonstrated for low-loss blue and UV photonics [50, 51, 52] and also employed in recent integrated trapped-ion experiments [33, 53]. Similar schemes can be implemented at longer wavelengths utilizing e.g. barium ions; and extension of schemes utilization carrier suppression for EIT cooling to ion species with nonzero nuclear spin, including with alternate beam and polarization configurations, is an interesting direction for further exploration [54]. Experiments towards realization of the proposed schemes with ${}^{40}\text{Ca}^{+}{}$ using foundry-fabricated trap devices incorporating both alumina and silicon nitride waveguides [22] are in progress.